A 6 pole, 3 – phase, 60 – cycle alternator has 12 slots per pole and four conductor per slot. The winding is 5/6 pitch. There are 2, 500, 000 maxwells (0. 025 weber) entering the armature from each north pole and this flux is sinusoidally distributed along the air gap. The armature coils are all connected in series. The winding is wye – connected. Determine the open circuit emf of the alternator. 


1
Expert's answer
2021-07-05T04:54:49-0400

"K_d=\frac{\sin{\frac{m\beta}{2}}}{m\sin{\frac{\beta}{2}}}\\
m=\frac{12}{3}=4\\
\beta=\frac{180\degree}{n}\\
n=12\\
\beta=\frac{180\degree}{12}\\
\beta=15\degree\\
K_d=\frac{\sin{\frac{4\times 15\degree}{2}}}{4\sin{\frac{15\degree}{2}}}=0.96\\
Z=4\times 6\times \frac{12}{3}=96\\
K_c=\frac{5}{6}=0.83\\
f=60Hz\\
\Phi=0.025Wb\\
E=2.22\times{{f\times{\Phi}}\times{Z}}\times{K_cK_d}\times \sqrt{3}\\
E=367.7V"